OpenStudy (anonymous):

Let u, v and w be three non-zero vectors no two of which are parallel. Show that if u × v = v × w = w × u then u + v + w = 0.

OpenStudy (zarkon):

lolok at \[(u+v+w) \times u\] \[(u+v+w) \times v\] \[(u+v+w) \times w\]

OpenStudy (zarkon):

do you see what you need to do?

OpenStudy (anonymous):

Yep. Thanks.